OpenStudy (anonymous):

a recipe calls for 3/6 cup of blackberries and 1/5 cup of rasberries .what is the total amount of berries needed for the recipe?

OpenStudy (anonymous):

You need to find a common denominator. Since 6 and 5 both go into 30... (3/6)=(15/30) and (1/5)=(6/30) therefore, you answer is (15/30)+(6/30)= 21/30 total cup of berries.

OpenStudy (anonymous):

which is also equal to 7/10 cup of berries